OpenStudy (anonymous):

A hotel has two ice machines in its kitchen. How many hours would it take both machines to make 60 pounds of ice? Round to the nearest tenth. 1st machine = 35 Ib/day 2nd machine= 25Ib/days

OpenStudy (redplanet2001):

Machine 1-41.1 hours Machine 2-57.6 hours

OpenStudy (anonymous):

Would you have to combined them

OpenStudy (redplanet2001):

In the case of it being combined it would be 24 hours but that seems to easy.

OpenStudy (redplanet2001):

And it wouldn't say round to the nearest tenth.

OpenStudy (anonymous):

Oh thanks

OpenStudy (redplanet2001):

Best response?

OpenStudy (anonymous):

Could you explain how you got it

OpenStudy (redplanet2001):

\[(60 \div 35) \times 24 \] and \[(60 \div 25) \times 24\]

OpenStudy (mertsj):

In 1 hour the first machine makes 35/24 pounds of ice In 1 hour the second machine makes 25/24 pounds of ice In x hours the first machine makes (35/24)x pounds of ice In x hours the second machine makes (25/24)x pounds of ice Together the two machines make 60 pounds of ice in x hours: \[\frac{35}{24}x+\frac{25}{24}x=60\]